a man bought 4 cups of coffee and left a $7 tip a woman bought 8 cups of coffee and left a $2 tip if they paid the same amount how much was each cup of coffee?
Answer by ankor@dixie-net.com(22740)

a man bought 4 cups of coffee and left a $7 tip a woman bought 8 cups of coffee and left a $2 tip if they paid the same amount how much was each cup of coffee?
:
let c = amt for each cup of coffee
woman amt = man amt
8c + 2 = 4c + 7
8c - 4c = 7 - 2
4c = 5
c = 5/4
c = $1.25